A sample of 400 observations will be taken from an infinite population. The population proportion equals 0.8. Find the probability that the sample proportion will be greater than 0.74.
Answer #1

Solution

Given that,

\sigma\hat p =  \sqrt[p( 1 - p ) / n] = \sqrt [(0.8 * 0.2) / 400 ] = 0.02

P( \hat p > 0.74) = 1 - P( \hat p < 0.74)

= 1 - P(( \hat p - \mu \hat p ) / \sigma \hat p < (0.74 - 0.8) / 0.02)

= 1 - P(z < -3)

= 1 - 0.0013

= 0.9987

probability = 0.9987